OpenStudy (anonymous):

Find the LCM 14a^6b^6, 2a^2b5, 21a^8b^2

OpenStudy (anonymous):

\[2 \times 3 \times 7 \times a^8 \times b^6 \] the LCM is below \[42a^8b^6\] that is assuming you ment b^5 when you wrote 2a^2b5